# Quantifying the Dunkelflaute — An analysis of variable renewable energy droughts in Europe

### v3

@@ 2.7 Geographical balancing inside the European renewable energy super drought
Figure 9 shows four snapshots of selected hours within the European portfolio super drought, as identified by the drought mass metric at the turn of the years 1996/97. Each panel shows the renewable drought severity present in the displayed hour for isolated countries in the island scenario (maps, top), illustrated by the lowest applying threshold. The corresponding drought severity in the European copperplate scenario, which is identical across all countries at each time step due to unconstrained geographical balancing, is illustrated as vertical bars at the bottom. While the super drought affects multiple European countries simultaneously, the severity varies across countries and over time. Comparing the island scenario and the European copperplate shows that renewable drought severity under unconstrained geographical balancing is always less pronounced than in the most affected isolated countries at the same hour. This is because the assumed European interconnection allows to mitigate energy shortages in one country by leveraging higher renewable availability in others. That is, geographical balancing is feasible even within the most severe pan-European drought event in the data. Further, even within the European super drought, there are hours with relatively high renewable availability in some countries (compare white-colored countries in Figures 9c and 9d). A complementary animated graph available as supplementary information [kittel_high_2025] shows the progression of the drought severity inside the European super drought, covering December 1996 and January 1997. The animation shows that the super drought comprises sequences of brief but more severe droughts, both in individual countries (different colors on maps on the left-hand side) and on the European level (different colors between maps on the right-hand side). It further shows that very severe droughts that cover most of Europe (e.g., also visible in Figure 9a) are relatively brief and occur rarely, even inside the super drought event. That is, some renewable generation potential is always available somewhere in Europe. Even during the most extreme droughts, this can balance low renewable availability across technologies and regions to some extent.
@@ 3 Discussion
Previous research has highlighted the importance of diversifying variable renewable energy supply [heide_seasonal_2010, jurasz_review_2020, harrison-atlas_temporal_2022] and advancing the integration of European power systems via transmission grid expansion [schaber_parametric_2012, fursch_role_2013, roth_geographical_2023] for realizing cost-efficient renewable energy systems. Based on our analysis, we not only underscore this notion but extend it by arguing that technology portfolios and geographical balancing are also important strategies for effectively dealing with rare and extreme renewable energy droughts. We also show that the most severe VRE portfolio droughts largely drive the need for long-duration storage in energy systems fully based on variable renewable energy sources.
@@
Our multi-threshold analysis facilitates a comprehensive characterization of renewable drought events in Europe, solely based on VRE availability time series. Our approach, and particularly the drought mass metric, is not prone to artifacts that may arise in drought analyses relying on energy system modeling such as the duration of the planning horizon or technology cost assumptions. These may substantially affect optimal long-duration storage investments [dowling_role_2020]. Yet, methodological challenges remain. Drought analyses purely based on renewable availability time series cannot take into account real-world interconnection levels, which are between our extreme assumptions in the island and copperplate scenarios. Drought characteristics of scenarios with cross-regional exchange are thus likely between these corner solutions. For instance, the duration of the European super drought is likely to last longer than the 55 days identified for the copperplate scenario, but shorter than identified here for isolated countries (Figure 8), depending on future interconnection capacities between countries.
@@
Further, the drought mass indicator does not consider a range of aspects that energy system models do. For instance, it does not capture periods with very high renewable availability that may arise between or after extreme droughts. Energy models with perfect foresight optimize long-duration storage operations considering not only droughts but also high-availability periods. This may result in major storage discharge periods that occur at different times or last longer than the most extreme drought periods identified by the drought mass indicator. In addition, the drought mass does not account for storage conversion losses or renewable curtailment. The latter may be required even within longer-lasting droughts in case of brief periods with very high availability that exceed storage charging capacities [ruhnau_storage_2022]. It is possible to integrate conversion losses into drought identification methods, but this comes at the cost of higher complexity [kittel_measuring_2024].
@@
Due to computational limits, energy system models are often solved for a limited number of weather years, or even for only a single weather year. While substantial inter-annual variation in optimal energy model outcomes has been documented in the literature [pfenninger_long-term_2016, staffell_using_2016, pfenninger_dealing_2017, zeyringer_designing_2018, collins_impacts_2018, schlott_impact_2018, kaspar_climatological_2019, hill_effects_2021, tong_geophysical_2021, grochowicz_intersecting_2023, gunn_spatial_2023, ruggles_planning_2024], our analysis shows that extreme portfolio droughts also vary substantially across different years and regions. This likely aggravates inter-annual variations of energy model outcomes, particularly in terms of long-duration system flexibility needs. Accordingly, it appears desirable to corroborate the findings and conclusions of previous model-based energy system studies that are based on a small number of weather years or even only a single one and do not include extreme VRE droughts. For example, the European Ten Year Network Development Plan 2022 draws on the single weather year 2009 [entso-e_tyndp_2022], while the newer Ten Year Network Development Plan 2024 uses the year 2019 [entso-e_tyndp_2024]. Long-term energy scenarios developed for the European Commission generally lack transparency about actual weather years used, but appear to be based on a very limited number of weather years [european_commission_directorate_general_for_energy_eu_2021]. The same holds true for influential climate neutrality scenarios for Germany [bdi_klimapfade_2021, fraunhofer_isi_langfristszenarien_2024, prognos_klimaneutrales_2021, luderer_deutschland_2021]. Our analysis shows that the winter of 1996/97 including the pan-European super drought is particularly relevant for weather-resilient planning of the European energy system. It further indicates relevant weather years for single-country analysis. For any model-based energy system analysis that aims to shed light on how to deal with extreme renewable energy droughts, we suggest a two-step approach: first, use a large weather data set and identify the weather year with the most extreme drought event, purely drawing on renewable time series analysis as showcased in this paper; and second, ensure to include this particular year when running complex, numerically constrained energy system models.
@@
Energy model planning horizons that are in line with a single calendar year further appear unsuitable for investigating weather-resilient future scenarios as extreme, storage-defining drought events may extend across the turn of years. Instead, we recommend adopting a planning horizon that captures relevant renewable seasonality patterns. For European settings, this may require at least bi-annual or single-year summer-to-summer planning horizons. Considering the typical annual cycles of European hydro reservoirs, which are not in the focus of our analysis, a spring-to-spring horizon might also be useful. Multi-annual planning horizons would be preferable but may come with increasing numerical challenges [grochowicz_using_2024, ruggles_planning_2024], and the length of the model period can impact optimal storage capacities [dowling_role_2020].
@@
Additionally, multi-sectoral and technology-rich energy system models with detailed grid representations often have to use aggregated temporal resolutions to reduce the computational burden [nahmmacher_carpe_2016, kittel_temporal_2022, teichgraeber_time-series_2022, goke_adequacy_2022]. Our results indicate that the dynamics of storage-defining extreme drought events should adequately be reflected by representative time slices used in such models [sanchez-perez_effect_2022]. This is even more relevant for integrated assessment models, which are also used for long-term energy system analyses [gong_bidirectional_2023, schreyer_distinct_2024], but usually come with even coarser time resolutions than energy system models. How to design respective temporal aggregation methods to properly represent extreme droughts that consist of a series of short extreme drought events during a prolonged period with low average renewable availability remains a methodological challenge.
@@
We see several promising avenues for future research. For example, a methodological comparison with alternative approaches that measure deviations from actual renewable availability from time-varying reference profiles [stoop_climatological_2024, antonini_identification_2024], would be of interest. We note that a wide range of different datasets of renewable availability are used for energy system modeling, with sometimes conflicting model outcomes and policy recommendations [kies_critical_2021]. Assessing the sensitivity of VRE drought characteristics to such datasets [kies_critical_2021, morin_reducing_2025] could provide valuable feedback to the meteorological research community that generates and continuously improves such datasets [craig_overcoming_2022]. Further, the temporal extent of most data sets is limited to a few decades. Statistically robust findings for low-probability VRE drought events with high impact are likely to require longer records of weather data, which also might involve a need for synthetic datasets, e.g., applied in Gangopadhyay et al. [gangopadhyay_role_2022]. Moreover, drought analyses for modeled future weather data that incorporate the effects of global climate change would be desirable [kapica_potential_2024, dolores_projections_2025]. Finally, using our methodology to characterize renewable energy droughts in other world regions and comparing these with our findings for Europe would be of interest. While parameters like the cut-off threshold of the drought mass metric ought to be reviewed, we argue that our multi-threshold framework is particularly useful for respective analyses in world regions where renewable availability is characterized by diurnal or seasonal variability, e.g., related to regional weather phenomena such as monsoon events.
@@ 4.1 Variable renewable energy drought definition and identification
Our VRE drought analysis is based on hourly VRE availability factor time series. An hourly availability factor, also referred to as capacity factor, ranges between 0 and 1 and indicates the generation potential (in MWh) of a VRE source in a specific region and hour normalized by its capacity (in MW). We use the variable-duration mean below threshold (VMBT) algorithm for drought identification as introduced and discussed in Kittel & Schill [kittel_measuring_2024]. This method searches for periods with a moving average below a given drought qualification threshold by iteratively decreasing the event duration. In each iteration, the algorithm sets the averaging interval to the respective event duration, starting with very long-lasting events and iteratively continuing to such that last only a few hours. A moving average below the drought threshold identifies a drought event. It is excluded from subsequent iterations, in which the averaging interval decreases further and additional (shorter) events are identified.
@@
This iterative procedure overcomes shortcomings of previous research [kittel_measuring_2024]. It pools adjacent periods that independently may not qualify as VRE drought, identifies unique events with the longest duration possible, avoids double counting as well as overlaps with adjacent events, and ensures that the full temporal extent of drought periods is captured [kittel_measuring_2024]. To cover the full spectrum of VRE droughts, the VMBT algorithm iteratively searches for periods ranging in duration from two years to one hour in descending order. The code and input data of the drought analysis tool are freely available on GitLab at https://gitlab.com/diw-evu/variable_renewable_energy_droughts_analyzer.
@@
While threshold-based renewable drought identification methods are commonly applied in different strands of the academic literature, previous threshold-based analyses of renewable droughts across multiple regions, technologies, or technology portfolios often apply uniform, exogenously defined thresholds e.g., [leahy_persistence_2013, cannon_using_2015, patlakas_low_2017, kaspar_climatological_2019, ohlendorf_frequency_2020, kies_critical_2021, potisomporn_extreme_2024, potisomporn_extreme_2024, abdelaziz_assessing_2024, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022]). Such approaches implicitly assume that one unit of generation capacity has a uniform generation potential across all region-technology settings. However, annual generation potentials may differ substantially between regions and technologies due to varying meteorological conditions, which makes comparisons of the resulting drought characteristics difficult. To address this, threshold-based drought analyses should account for cross-regional and cross-technological variations in renewable generation potential when defining drought thresholds. Thresholds scaled relative to the global mean availability of each setting (termed here as “relative thresholds”), reflect these variations [kittel_measuring_2024].
@@
Additionally, as highlighted in the literature review, no consensus has yet been established regarding the selection of concrete threshold values. On the contrary, they often appear arbitrarily chosen or lack clear justification [kittel_measuring_2024].
@@ 4.3 Power sector modeling to determine long-duration storage use
We use a stylized version of the open-source power sector model Dispatch and Investment Evaluation Tool with Endogenous Renewables (DIETER) to analyze the interaction between VRE droughts and long-duration storage needs in a fully renewable European power sector. DIETER is a linear optimization model that determines least-cost capacity and dispatch decisions based on an hourly resolution [zerrahn2017, gaete-morales_dieterpy_2021]. Different versions of the model have been used to study various aspects of VRE integration and their interaction with other flexibility options or sector coupling technologies [zerrahn2018, schill_long-run_2018, schill_electricity_2020, schill_flexible_2020, kittel_renewable_2022, roth_geographical_2023, kirchem_power_2023, gong_bidirectional_2023, gaete-morales_power_2024, roth_power_2024]. Here, we use a model version that includes 33 European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans). As very long-lasting VRE droughts may span across the turn of a calendar year, we extend DIETER’s planning horizon to two full years. For transparency and reproducibility, we provide the model code and a manual in a public repository under permissive licenses available on GitLab at https://gitlab.com/diw-evu/projects/quantifying_the_dunkelflaute_energy_system_analysis.
The model features green hydrogen technologies, covering its generation, storage, and reconversion to electricity. We assume that hydrogen cavern storage can be expanded without restrictions in every country. This is a deliberate simplification, as we aim to illustrate the relation between VRE portfolio droughts and long-duration electricity storage, but not to derive policy-relevant geographical allocations of hydrogen storage across Europe [caglayan_technical_2020, talukdar_techno-economic_2024]. We focus on hydrogen used for long-duration electricity storage and abstract from additional hydrogen demand of other sectors, hydrogen imports from other world regions, and hydrogen exchange between countries.
@@ 4.4 Input data
For the renewable drought analysis, we use country-level VRE availability time series from the Pan-European Climate Database provided by the European Network of Transmission System Operators for Electricity [de_felice_entso-e_2022], comprising 38 weather years from 1982 to 2019. These data have been derived from reanalysis data, which has been converted to renewable availability by the Transmission System Operators. The database has been used for policy-relevant strategic reports, such as the Ten Year Network Development Plan (TYNDP) 2022 or the European Resource Adequacy Assessment 2021. We use exogenous renewable capacity assumptions to generate the capacity-weighted VRE portfolio time series as explained in the next paragraph. These assumptions stem from the TYNDP 2022 (scenario Distributed Energy), which optimized these portfolios for a climate-neutral European power sector in 2050. For Germany, we update the capacity targets according to the latest government targets for 2045 [bundesamt_fur_justiz_gesetz_2024], in which Germany aims to achieve climate-neutrality.

### v4

@@ 2.7 Geographical balancing inside the European renewable energy super drought
Figure 9 shows four snapshots of selected hours within the European portfolio super drought, as identified by the drought mass metric at the turn of the years 1996/97. Each panel shows the renewable drought severity present in the displayed hour for isolated countries in the island scenario (maps, top), illustrated by the lowest applying threshold. The corresponding drought severity in the European copperplate scenario, which is identical across all countries at each time step due to unconstrained geographical balancing, is illustrated as vertical bars at the bottom. While the super drought affects multiple European countries simultaneously, the severity varies across countries and over time. Comparing the island scenario and the European copperplate shows that renewable drought severity under unconstrained geographical balancing is always less pronounced than in the most affected isolated countries at the same hour. This is because the assumed European interconnection allows to mitigate energy shortages in one country by leveraging higher renewable availability in others. That is, geographical balancing is feasible even within the most severe pan-European drought event in the data. Further, even within the European super drought, there are hours with relatively high renewable availability in some countries (compare white-colored countries in Figures 9c and 9d). Supplementary Movie 1 provides a complementary animated graph, which shows the progression of the drought severity inside the European super drought, covering December 1996 and January 1997. The animation shows that the super drought comprises sequences of brief but more severe droughts, both in individual countries (different colors on maps on the left-hand side) and on the European level (different colors between maps on the right-hand side). It further shows that very severe droughts that cover most of Europe (e.g., also visible in Figure 9a) are relatively brief and occur rarely, even inside the super drought event. That is, some renewable generation potential is always available somewhere in Europe. Even during the most extreme droughts, this can balance low renewable availability across technologies and regions to some extent.
@@ 3 Discussion
Previous research has highlighted the importance of diversifying variable renewable energy supply [37, 44, 35] and advancing the integration of European power systems via transmission grid expansion [81, 26, 77] for realizing cost-efficient renewable energy systems. Based on our analysis, we not only underscore this notion but extend it by arguing that technology portfolios and geographical balancing are also important strategies for effectively dealing with rare and extreme renewable energy droughts. We also show that the most severe VRE portfolio droughts largely drive the need for long-duration storage in energy systems fully based on variable renewable energy sources.
@@
Our multi-threshold analysis facilitates a comprehensive characterization of renewable drought events in Europe, solely based on VRE availability time series. Our approach, and particularly the drought mass metric, is not prone to artifacts that may arise in drought analyses relying on energy system modeling such as the duration of the planning horizon or technology cost assumptions. These may substantially affect optimal long-duration storage investments [21]. Yet, methodological challenges remain. Drought analyses purely based on renewable availability time series cannot take into account real-world interconnection levels, which are between our extreme assumptions in the island and copperplate scenarios. Drought characteristics of scenarios with cross-regional exchange are thus likely between these corner solutions. For instance, the duration of the European super drought is likely to last longer than the 55 days identified for the copperplate scenario, but shorter than identified here for isolated countries (Figure 8), depending on future interconnection capacities between countries.
@@
Further, the drought mass indicator does not consider a range of aspects that energy system models do. For instance, it does not capture periods with very high renewable availability that may arise between or after extreme droughts. Energy models with perfect foresight optimize long-duration storage operations considering not only droughts but also high-availability periods. This may result in major storage discharge periods that occur at different times or last longer than the most extreme drought periods identified by the drought mass indicator. In addition, the drought mass does not account for storage conversion losses or renewable curtailment. The latter may be required even within longer-lasting droughts in case of brief periods with very high availability that exceed storage charging capacities [79]. It is possible to integrate conversion losses into drought identification methods, but this comes at the cost of higher complexity [51].
@@
Due to computational limits, energy system models are often solved for a limited number of weather years, or even for only a single weather year. While substantial inter-annual variation in optimal energy model outcomes has been documented in the literature [65, 90, 66, 103, 16, 86, 46, 38, 94, 32, 34, 78], our analysis shows that extreme portfolio droughts also vary substantially across different years and regions. This likely aggravates inter-annual variations of energy model outcomes, particularly in terms of long-duration system flexibility needs. Accordingly, it appears desirable to corroborate the findings and conclusions of previous model-based energy system studies that are based on a small number of weather years or even only a single one and do not include extreme VRE droughts. For example, the European Ten Year Network Development Plan 2022 draws on the single weather year 2009 [23], while the newer Ten Year Network Development Plan 2024 uses the year 2019 [22]. Long-term energy scenarios developed for the European Commission generally lack transparency about actual weather years used, but appear to be based on a very limited number of weather years [24]. The same holds true for influential climate neutrality scenarios for Germany [5, 25, 71, 56]. Our analysis shows that the winter of 1996/97 including the pan-European super drought is particularly relevant for weather-resilient planning of the European energy system. It further indicates relevant weather years for single-country analysis. For any model-based energy system analysis that aims to shed light on how to deal with extreme renewable energy droughts, we suggest a two-step approach: first, use a large weather data set and identify the weather year with the most extreme drought event, purely drawing on renewable time series analysis as showcased in this paper; and second, ensure to include this particular year when running complex, numerically constrained energy system models.
@@
Energy model planning horizons that are in line with a single calendar year further appear unsuitable for investigating weather-resilient future scenarios as extreme, storage-defining drought events may extend across the turn of years. Instead, we recommend adopting a planning horizon that captures relevant renewable seasonality patterns. For European settings, this may require at least bi-annual or single-year summer-to-summer planning horizons. Considering the typical annual cycles of European hydro reservoirs, which are not in the focus of our analysis, a spring-to-spring horizon might also be useful. Multi-annual planning horizons would be preferable but may come with increasing numerical challenges [33, 78], and the length of the model period can impact optimal storage capacities [21].
@@
Additionally, multi-sectoral and technology-rich energy system models with detailed grid representations often have to use aggregated temporal resolutions to reduce the computational burden [60, 49, 93, 30]. Our results indicate that the dynamics of storage-defining extreme drought events should adequately be reflected by representative time slices used in such models [80]. This is even more relevant for integrated assessment models, which are also used for long-term energy system analyses [31, 87], but usually come with even coarser time resolutions than energy system models. How to design respective temporal aggregation methods to properly represent extreme droughts that consist of a series of short extreme drought events during a prolonged period with low average renewable availability remains a methodological challenge.
@@
We see several promising avenues for future research. For example, a methodological comparison with alternative approaches that measure deviations from actual renewable availability from time-varying reference profiles [91, 3], would be of interest. We note that a wide range of different datasets of renewable availability are used for energy system modeling, with sometimes conflicting model outcomes and policy recommendations [47]. Assessing the sensitivity of VRE drought characteristics to such datasets [47, 59] could provide valuable feedback to the meteorological research community that generates and continuously improves such datasets [17]. Further, the temporal extent of most data sets is limited to a few decades. Statistically robust findings for low-probability VRE drought events with high impact are likely to require longer records of weather data, which also might involve a need for synthetic datasets, e.g., applied in Gangopadhyay et al. [29]. Moreover, drought analyses for modeled future weather data that incorporate the effects of global climate change would be desirable [45, 20]. Finally, using our methodology to characterize renewable energy droughts in other world regions and comparing these with our findings for Europe would be of interest. While parameters like the cut-off threshold of the drought mass metric ought to be reviewed, we argue that our multi-threshold framework is particularly useful for respective analyses in world regions where renewable availability is characterized by diurnal or seasonal variability, e.g., related to regional weather phenomena such as monsoon events.
@@ 4.1 Variable renewable energy drought definition and identification
Our VRE drought analysis is based on hourly VRE availability factor time series. An hourly availability factor, also referred to as capacity factor, ranges between 0 and 1 and indicates the generation potential (in MWh) of a VRE source in a specific region and hour normalized by its capacity (in MW). We use the variable-duration mean below threshold (VMBT) algorithm for drought identification as introduced and discussed in Kittel & Schill [51]. This method searches for periods with a moving average below a given drought qualification threshold by iteratively decreasing the event duration. In each iteration, the algorithm sets the averaging interval to the respective event duration, starting with very long-lasting events and iteratively continuing to such that last only a few hours. A moving average below the drought threshold identifies a drought event. It is excluded from subsequent iterations, in which the averaging interval decreases further and additional (shorter) events are identified.
@@
This iterative procedure overcomes shortcomings of previous research [51]. It pools adjacent periods that independently may not qualify as VRE drought, identifies unique events with the longest duration possible, avoids double counting as well as overlaps with adjacent events, and ensures that the full temporal extent of drought periods is captured [51]. To cover the full spectrum of VRE droughts, the VMBT algorithm iteratively searches for periods ranging in duration from two years to one hour in descending order.
@@
While threshold-based renewable drought identification methods are commonly applied in different strands of the academic literature, previous threshold-based analyses of renewable droughts across multiple regions, technologies, or technology portfolios often apply uniform, exogenously defined thresholds e.g., [53, 14, 64, 46, 62, 47, 69, 69, 1, 58, 57, 61]). Such approaches implicitly assume that one unit of generation capacity has a uniform generation potential across all region-technology settings. However, annual generation potentials may differ substantially between regions and technologies due to varying meteorological conditions, which makes comparisons of the resulting drought characteristics difficult. To address this, threshold-based drought analyses should account for cross-regional and cross-technological variations in renewable generation potential when defining drought thresholds. Thresholds scaled relative to the global mean availability of each setting (termed here as “relative thresholds”), reflect these variations [51].
@@
Additionally, as highlighted in the literature review, no consensus has yet been established regarding the selection of concrete threshold values. On the contrary, they often appear arbitrarily chosen or lack clear justification [51].
@@ 4.3 Power sector modeling to determine long-duration storage use
We use a stylized version of the open-source power sector model Dispatch and Investment Evaluation Tool with Endogenous Renewables (DIETER) to analyze the interaction between VRE droughts and long-duration storage needs in a fully renewable European power sector. DIETER is a linear optimization model that determines least-cost capacity and dispatch decisions based on an hourly resolution [102, 28]. Different versions of the model have been used to study various aspects of VRE integration and their interaction with other flexibility options or sector coupling technologies [101, 82, 84, 83, 50, 77, 48, 31, 27, 76]. Here, we use a model version that includes 33 European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans). As very long-lasting VRE droughts may span across the turn of a calendar year, we extend DIETER’s planning horizon to two full years.
The model features green hydrogen technologies, covering its generation, storage, and reconversion to electricity. We assume that hydrogen cavern storage can be expanded without restrictions in every country. This is a deliberate simplification, as we aim to illustrate the relation between VRE portfolio droughts and long-duration electricity storage, but not to derive policy-relevant geographical allocations of hydrogen storage across Europe [12, 92]. We focus on hydrogen used for long-duration electricity storage and abstract from additional hydrogen demand of other sectors, hydrogen imports from other world regions, and hydrogen exchange between countries.
@@ 4.4 Input data
For the renewable drought analysis, we use country-level VRE availability time series from the Pan-European Climate Database provided by the European Network of Transmission System Operators for Electricity [18], comprising 38 weather years from 1982 to 2019. These data have been derived from reanalysis data, which has been converted to renewable availability by the Transmission System Operators. The database has been used for policy-relevant strategic reports, such as the Ten Year Network Development Plan (TYNDP) 2022 or the European Resource Adequacy Assessment 2021. We use exogenous renewable capacity assumptions to generate the capacity-weighted VRE portfolio time series as explained in the next paragraph. These assumptions stem from the TYNDP 2022 (scenario Distributed Energy), which optimized these portfolios for a climate-neutral European power sector in 2050. For Germany, we update the capacity targets according to the latest government targets for 2045 [11], in which Germany aims to achieve climate-neutrality.